OpenStudy (anonymous):

what is the limit as x approaches infinity of the square root of (9x^2+x)-3x

OpenStudy (anonymous):

Do you mean: \(\lim_{x \rightarrow \infty} \sqrt{9x^2 + x} - 3x \)?

OpenStudy (anonymous):

If so, try multiplying the limit by: \[ \frac{3x + \sqrt{9x^2 + x}}{3x + \sqrt{9x^2 + x}} \]That is, rationalize the expression first.

OpenStudy (dumbcow):

then after that, divide top/bottom by x remember x = sqrt(x^2)